Write a cubic function that passes through the following points: (-2,0) (3,0) (-1,0) and (1,2). Explain how you got the answer
Answer by TimothyLamb(4379)

the general standard cubic:
ax^3 + bx^2 + cx + d = 0
---
from the first 3 points (-2,0) (3,0) (-1,0), we have 3 roots at:
(x + 2)(x - 3)(x + 1)
---
so the equation is as follows, where k is an unknown constant:
(x + 2)(x - 3)(x + 1) + k = 0
(xx - x - 6)(x + 1) + k = 0
xxx + xx - xx - x - 6x - 6 + k = 0
xxx - 7x - 6 + k = 0
---
but from the 4th point (1,2) we know that:
xxx - 7x - 6 + k = 0
1*1*1 - 7*1 - 6 + k = 0
-12 + k = 0
k = 12
---
combining the constants, the equation is:
xxx - 7x - 6 + k = 0
xxx - 7x - 6 + 12 = 0
xxx - 7x + 6 = 0
---
answer:
x^3 - 7x + 6 = 0
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---